# CamSAM2: Segment Anything Accurately in Camouflaged Videos

### v2

@@ 3.3 Implicit Object-aware Fusion
Early-layer features from the image encoder capture high-resolution details, such as edges and textures, essential for distinguishing subtle differences between the camouflaged object and the background. In contrast, deeper layers focus on high-level semantic information. In SAM2, memory-conditioned features are computed by only conditioning high-level semantic features on the memory, without using detailed features from early layers. These early-layer features are implicit object-aware, as features for background and non-relevant objects also exist with similar magnitude. To this end, we propose an IOF module that fuses these implicit object-aware features with memory-conditioned features.
For SAM2, three feature maps from the Hiera image encoder [41] are extracted for each frame, i.e., $L=3$. We have $\mathbf{F}_{t}^{0}$, $\mathbf{F}_{t}^{1}$, and $\mathbf{F}_{t}^{2}$ for the current frame $\mathbf{I}_{t}$. We denote the memory-conditioned feature as $\mathbf{F}_{t}^{mem}$, encoded by the memory-attention module on $\mathbf{F}_{t}^{2}$, as in [39]. The high-resolution features $\mathbf{F}_{t}^{0}$ and $\mathbf{F}_{t}^{1}$ are fused with $\mathbf{F}_{t}^{mem}$ via compression modules and point-wise addition to create a refined feature representation $\mathbf{F}_{t}^{iof}\in\mathbb{R}^{c_{0}\times h_{0}\times w_{0}}$, where a compression module $C(\cdot)$ consists of two convolutional layers, followed by an upsampling layer, to create compact representations. This process is given by:
@@ 3.4 Explicit Object-aware Fusion
After obtaining $\mathbf{F}_{t}^{iof}$, we further refine it by EOF, which exploits explicit object-aware information from the current frame and previous frames, through employing object mask logits and object prototypes (see §3.5). We have three steps to fuse informative features. First, feature $\mathbf{F}_{t}^{{iof}}$, with shape $\mathbb{R}^{c_{0}\times h_{0}\times w_{0}}$ from the previous IOF module, is directly concatenated with SAM2’s mask logits $\mathbf{R}_{t}$, which has shape $\mathbb{R}^{1\times h_{0}\times w_{0}}$. This concatenated feature is then processed through a convolutional layer to reduce the channels back to $c_{0}$, resulting in output with the original shape $\mathbb{R}^{c_{0}\times h_{0}\times w_{0}}$, denoted as:
@@
Second, $\mathbf{F}_{t}^{eof}$ goes through a cross-attention layer. Prototypes generated from previous frames, representing clustered camouflaged features, serve as informative priors to help distinguish the camouflaged object from its background. A cross-attention mechanism takes $\mathbf{F}_{t}^{eof}$ as a query, and leverages these prototypes as keys and values, effectively exploiting the information within the object prototypes to refine $\mathbf{F}_{t}^{eof}$. Formally, we update $\mathbf{F}_{t}^{eof}$ by conducting cross-attention with prototypes $\mathcal{P}_{t}=\{\mathbf{P}_{0},\mathbf{P}_{1},\ldots,\mathbf{P}_{t-1}\}$ from previous frames, given by:
@@
Finally, we calculate the mask logits $\mathbf{R}_{t}^{c}$ of CamSAM2, by processing the output decamouflaged token $\mathbf{T}^{\prime}$ through an MLP layer, and then performing point-wise product with the $\mathbf{F}_{t}^{eof\prime}$, as shown below:
@@ 3.5 Object Prototype Generation
To effectively represent the camouflaged features within the mask (object) region, we employ Farthest Point Sampling (FPS) [31] to identify $k$ points within the predicted mask region, which act as cluster centers. This approach ensures that the sampled points are well-distributed throughout the mask, capturing diverse and important characteristics of the camouflaged object. Then, we group all pixels in the predicted mask region into $k$ clusters by conducting one-iteration k-means, using the sampled $k$ points as initial centers. The prototype of each cluster is represented as the mean of the spatial features of the points in the cluster. This prototype generation process is denoted as $\mathcal{F}_{p}$, as shown in:
@@
3.6 Training and Inference Strategies
Training Strategies.  We simulate interactive prompting of the model in the training process, prompting on the first frame of the sampled sequence. Following the training strategy of SAM2, we use three types of prompts (mask, bounding box, 1-click point of foreground) for training, with the probability of 0.5, 0.25, and 0.25, respectively.
@@ 3.6 Training and Inference Strategies
Inference.  During inference, we provide a prompt at the first frame of a video, following [29]. Our final output is the average of the logits of SAM2 and CamSAM2 masks for the error correction.
@@
EGNet [60] ResNet-50 111.7 - 54.7 11.0 3.5 13.6 57.4 14.3 9.6
BASNet [37] ResNet-50 87.1 - 56.1 15.4 4.2 17.3 59.8 19.0 13.7
CPD [48] ResNet-50 47.9 - 56.1 12.1 4.1 15.2 61.3 16.2 11.3
PraNet [11] ResNet-50 32.6 - 61.4 26.6 3.0 29.6 67.4 31.1 23.4
SINet [10] ResNet-50 48.9 - 59.8 23.1 2.8 25.6 69.9 27.7 20.2
SINet-V2 [12] Res2Net-50 27.0 - 58.8 20.4 3.1 22.9 64.2 24.5 18.0
PNS-Net [16] ResNet-50 142.9 - 52.6 5.9 3.5 8.4 53.0 8.4 5.4
RCRNet [52] ResNet-50 53.8 - 55.5 13.8 3.3 15.9 52.7 17.1 11.6
MG [53] VGG 4.8 - 53.0 16.8 6.7 19.5 56.1 18.1 12.7
SLT-Net-LT [7] PVTv2-B5 82.3 - 63.1 31.1 2.7 33.1 75.9 36.0 27.2
ZoomNeXt [33] PVTv2-B5 84.8 - 73.4 47.6 1.0 49.7 73.6 49.7 42.2
SAM2 [39] Hiera-T 38.9 1-click 68.2 50.7 7.7 52.5 73.6 52.1 44.8
@@
SAM-PM [29] ViT-L 303.0 mask 72.8 56.7 0.9 - 81.3 59.4 50.2
SAM2 [39] Hiera-T 38.9 mask 84.7 76.0 0.4 76.9 91.9 77.1 67.9
@@
Table 1: Comparisons between our method and existing approaches on MoCA-Mask. CamSAM2 outperforms the existing method by achieving new state-of-the-art performance. The results of all these methods (excluding SAM2) are from corresponding publications. The best results are shown in bold. $\uparrow$: the higher the better, $\downarrow$: the lower the better.
@@
Figure 3: Qualitative comparisons between SAM2 and CamSAM2 using 1-click prompt with the Hiera-T backbone on two MoCA-Mask clips. From top to bottom: the input frames, SAM2’s results, CamSAM2’s results, and ground-truth masks. CamSAM2 demonstrates improved accuracy in segmenting camouflaged objects, especially in complex backgrounds, as shown by the circles. Best viewed in color.

### v3

@@ 3.3 Implicit object-aware fusion
Early-layer features from the image encoder capture high-resolution details, such as edges and textures, essential for distinguishing subtle differences between the camouflaged object and the background. These early-layer features are implicit object-aware, as features for background and non-relevant objects also exist with similar magnitude. In contrast, deeper layers focus on high-level semantic information. In SAM2, memory-conditioned features are computed by only conditioning high-level semantic features, without using detailed features from early layers. To this end, we propose an IOF module to fuse these implicit object-aware features with memory-conditioned features.
For SAM2, three feature maps from the Hiera image encoder [ryali2023hiera] are extracted for each frame, i.e., $L=3$. We have $\mathbf{F}_{t}^{0}$, $\mathbf{F}_{t}^{1}$, and $\mathbf{F}_{t}^{2}$ for the current frame $\mathbf{I}_{t}$. We denote the memory-conditioned feature as $\mathbf{F}_{t}^{mem}$, encoded by the memory-attention module on $\mathbf{F}_{t}^{2}$, as in [ravi2025sam]. The high-resolution features $\mathbf{F}_{t}^{0}$ and $\mathbf{F}_{t}^{1}$ are fused with $\mathbf{F}_{t}^{mem}$ via compression modules and point-wise addition to create a refined feature representation $\mathbf{F}_{t}^{iof}\in\mathbb{R}^{c_{0}\times h_{0}\times w_{0}}$, where a compression module $C(\cdot)$ consists of two convolutional layers, followed by an upsampling layer. This process is given by:
@@ 3.4 Explicit object-aware fusion
After obtaining $\mathbf{F}_{t}^{iof}$, we further refine it by EOF, which exploits explicit object-aware information from the current frame and previous frames, through employing object mask logits and object prototypes (see §3.5). We have three steps to fuse informative features. First, feature $\mathbf{F}_{t}^{{iof}}$, with shape $\mathbb{R}^{c_{0}\times h_{0}\times w_{0}}$, is directly concatenated with SAM2’s mask logits $\mathbf{R}_{t}$, which has shape $\mathbb{R}^{1\times h_{0}\times w_{0}}$. This concatenated feature is then processed through a convolutional layer to reduce the channels back to $c_{0}$, resulting in the output with the original shape $\mathbb{R}^{c_{0}\times h_{0}\times w_{0}}$, denoted as:
@@
Second, $\mathbf{F}_{t}^{eof}$ goes through a cross-attention layer. Prototypes generated from previous frames, representing clustered camouflaged features, serve as informative priors to help distinguish the camouflaged object from its background. A cross-attention mechanism takes $\mathbf{F}_{t}^{eof}$ as a query, and leverages these prototypes as keys and values, effectively exploiting the information within the object prototypes to refine $\mathbf{F}_{t}^{eof}$. This design naturally suppresses outdated or irrelevant prototypes: if a prototype is inconsistent or suboptimal (e.g., due to occlusion or missing), it receives lower attention and contributes less to the output. This cross-attention mechanism provides robustness and avoids overfitting to specific regions. Formally, we update $\mathbf{F}_{t}^{eof}$ by conducting cross-attention with prototypes $\mathcal{P}_{t}=\{\mathbf{P}_{0},\mathbf{P}_{1},\ldots,\mathbf{P}_{t-1}\}$ from previous frames, given by:
@@
Finally, we calculate the mask logits $\mathbf{R}_{t}^{c}$ of CamSAM2, by processing the output decamouflaged token $\mathbf{T}^{\prime}$ through an MLP layer, then performing point-wise product with the $\mathbf{F}_{t}^{eof\prime}$, as shown below:
@@ 3.5 Object prototype generation
To effectively represent the camouflaged features within the mask (object) region, we employ Farthest Point Sampling (FPS) [moenning2003fast] to identify $k$ points within the predicted mask region, which act as cluster centers. This approach ensures that the sampled points are well-distributed throughout the mask, capturing diverse and important characteristics of the camouflaged object. Then, we group all pixels in the predicted mask region into $k$ clusters by conducting one-iteration k-means, using the sampled $k$ points as initial centers. The prototype of each cluster is represented as the mean of the spatial features of the points in the cluster. This prototype generation process is denoted as $\mathcal{F}_{p}$, as shown in:
@@
4 Experiments
4.1 Experimental setup
Datasets.
Our experiments
are conducted on three video datasets: two popular camouflaged animal datasets, MoCA-Mask [cheng2022implicit] and CAD [bideau2016s], and one camouflaged medical dataset, SUN-SEG [ji2022video]. The pioneering Moving Camouflaged Animals dataset (MoCA) [lamdouar2020betrayed] comprises 37K frames from 141 YouTube video sequences. The dataset MoCA-Mask is reorganized from the MoCA, containing 71 video sequences with 19,313 frames for training and 16 video sequences with 3,626 frames for testing, respectively, with pixel-wise ground-truth masks on every five frames. It also generates a MoCA-Mask-Pseudo dataset, which contains pseudo masks for unlabeled frames with a bidirectional optical-flow-based consistency check strategy. The Camouflaged Animal Dataset (CAD) includes 9 short videos in total that have 181 hand-labeled masks on every five frames. SUN-SEG is the largest benchmark for video polyp segmentation, derived from SUN-database [misawa2021development]. It consists of a training set with 112 clips (19,544 frames) and two test sets: SUN-SEG-Easy, containing 119 clips (17,070 frames), and SUN-SEG-Hard, comprising 54 clips (12,522 frames).
Training and inference.
We simulate interactive prompting of the model in the training process, prompting on the first frame of the sampled sequence. Following the training strategy of SAM2, we use three types of prompts (mask, bounding box, 1-click point of foreground) for training, with the probabilities of 0.5, 0.25, and 0.25, respectively.
@@ 4.1 Experimental setup
During inference, we provide a prompt at the first frame of a video, following [ravi2025sam, meeran2024sam]. Our final output is the average of the logits of SAM2 and CamSAM2 masks for the error correction. For more training and inferencing details, please see Appendix A.5.
Implementation details.
The proposed CamSAM2 is implemented with PyTorch [paszke2019pytorch]. CamSAM2 is initialized with the parameters of SAM2. We freeze all parameters used in SAM2 and initialize other parameters randomly. We set betas=(0.9,0.999)betas=(0.9,~0.999) for the optimizer Adam and use the learning rate of 1e-3. We train CamSAM2 on 4 NVIDIA RTX 4090 GPUs for 10 epochs. For camouflaged animal segmentation, we train the model using the MoCA-Mask-Pseudo training set and evaluate it on the MoCA-Mask test set and CAD. During inference, we apply the 1-click, box, and mask prompts only on the first frame of each video. For camouflaged polyp segmentation, we train the model using the SUN-SEG training set and perform inference using the mask prompt on the first frame of each video on the SUN-SEG-Easy and SUN-SEG-Hard test sets.
Evaluation metrics.
We adopt seven evaluation metrics to measure the quality of predicted pixel-wise masks: S-measure ($S_{m}$[Smeasure], F-measure ($F_{\beta}$[Fmeasure], weighted F-measure ($F_{\beta}^{\omega}$[wFmeasure], mean absolute error (MAE) [MAE], E-measure ($E_{m}$[Emeasure], mean Dice (mDice), and mean IoU (mIoU).
4.2 Experimental results
Table 1: Comparisons between our method and existing approaches on MoCA-Mask. CamSAM2 outperforms the existing method by achieving new state-of-the-art performance. “SAM2-FT” refers to a version of SAM2 in which the mask decoder is fine-tuned. The results of all these methods (excluding SAM2) are from the corresponding publications. The best results are shown in bold. $\uparrow$: the higher the better, $\downarrow$: the lower the better.
@@ 4.2 Experimental results
EGNet2019 [zhao2019egnet] ResNet-50 111.7 - 54.7 11.0 3.5 13.6 57.4 14.3 9.6
BASNet2019 [qin2019basnet] ResNet-50 87.1 - 56.1 15.4 4.2 17.3 59.8 19.0 13.7
CPD2019 [wu2019cascaded] ResNet-50 47.9 - 56.1 12.1 4.1 15.2 61.3 16.2 11.3
PraNet2020 [fan2020pranet] ResNet-50 32.6 - 61.4 26.6 3.0 29.6 67.4 31.1 23.4
SINet2020 [fan2020camouflaged] ResNet-50 48.9 - 59.8 23.1 2.8 25.6 69.9 27.7 20.2
SINet-V22021 [fan2021concealed] Res2Net-50 27.0 - 58.8 20.4 3.1 22.9 64.2 24.5 18.0
PNS-Net2021 [ji2021progressively] ResNet-50 142.9 - 52.6 5.9 3.5 8.4 53.0 8.4 5.4
RCRNet2019 [yan2019semi] ResNet-50 53.8 - 55.5 13.8 3.3 15.9 52.7 17.1 11.6
MG2021 [yang2021self] VGG 4.8 - 53.0 16.8 6.7 19.5 56.1 18.1 12.7
SLT-Net-LT2022 [cheng2022implicit] PVTv2-B5 82.3 - 63.1 31.1 2.7 33.1 75.9 36.0 27.2
ZoomNeXt2024 [pang2024zoomnext] PVTv2-B5 84.8 - 73.4 47.6 1.0 49.7 73.6 49.7 42.2
SAM22024 [ravi2025sam] Hiera-T 38.9 1-click 68.2 50.7 7.7 52.5 73.6 52.1 44.8
SAM2-FT Hiera-T 38.9 1-click 70.2 54.5 7.2 56.0 76.8 55.2 48.1
@@
SAM-PM2024 [meeran2024sam] ViT-L 303.0 mask 72.8 56.7 0.9 - 81.3 59.4 50.2
SAM22024 [ravi2025sam] Hiera-T 38.9 mask 84.7 76.0 0.4 76.9 91.9 77.1 67.9
SAM2-FT Hiera-T 38.9 mask 84.7 76.2 0.4 77.1 91.9 77.2 68.0
@@
Table 2: Detailed comparisons between SAM2 and CamSAM2 on MoCA-Mask. CamSAM2 consistently outperforms SAM2 for all considered prompt types and backbones. Improvements of CamSAM2 over SAM2 are shown in dark green.
       Model       Prompt       Hiera-T       Hiera-S
      mDice $\uparrow$        mIoU $\uparrow$       mDice $\uparrow$       mIoU $\uparrow$
SAM2 1-click 52.1 44.8 54.9 46.7
CamSAM2 64.3
(+12.2)
54.6
(+9.8)
68.0
(+13.1)
58.8
(+12.1)
SAM2 box 72.7 62.3 73.7 63.8
CamSAM2 75.5
(+2.8)
64.8
(+2.5)
76.4
(+2.7)
66.1
(+2.3)
SAM2 mask 77.1 67.9 80.3 70.7
CamSAM2 80.2
(+3.1)
70.5
(+2.6)
81.4
(+1.1)
71.7
(+1.0)
Results on MoCA-Mask.
Tab.˜1 compares four promptable methods on MoCA-Mask. For a fair comparison, we fine-tune the mask decoder (4.2M parameters) of SAM2 on MoCA-Mask. CamSAM2 clearly outperforms SAM-PM, SAM2, and the fine-tuned SAM2. Even with the 1-click prompt, CamSAM2 still outperforms SAM-PM, which uses the mask prompt. The promptable methods clearly outperform other non-promptable models. Tab.˜2 further compares SAM2 and CamSAM2 with the Hiera-T and Hiera-S backbones. CamSAM2 consistently outperforms SAM2 across all prompt types. With the 1-click prompt, CamSAM2 achieves mDice/mIoU gains of 12.2/9.8 with Hiera-T backbone and 13.1/12.1 with Hiera-S backbone, and further improves mIoU by 2.5/2.3 (box) and 2.6/1.0 (mask) over SAM2 on Hiera-T and Hiera-S backbones, respectively.
@@
Figure 3: Qualitative comparisons between SAM2 and CamSAM2 using 1-click prompt with the Hiera-T backbone on two MoCA-Mask clips. From top to bottom: the input frames, SAM2’s results, CamSAM2’s results, and ground-truth masks. CamSAM2 demonstrates improved accuracy in VCOS, especially in complex backgrounds, as shown by the circles. Best viewed in color.
Figure 4: Attention map visualization from SAM2 and CamSAM2 using point prompts with the Hiera-T backbone. From top to bottom: input frames, attention with SAM2 mask token, attention with decamouflaged token, and ground-truth masks. The higher attention regions are indicated by warmer colors.
Fig.˜3 presents qualitative comparisons on two MoCA-Mask video clips using 1-click prompts with Hiera-T as the backbone. Fig.˜4 shows attention maps from the last token-to-image cross-attention layer in the mask decoder, where the SAM2 output token or the decamouflaged token acts as the query and the image embedding as key and value. Compared to SAM2, CamSAM2 produces more focused and expansive attention around target objects, highlighting its improved spatial awareness and the effectiveness of our design for VCOS. Despite CamSAM2 introducing only a marginal parameter increase of 0.5M, it delivers substantial improvements while keeping all SAM2 parameters frozen, fully preserving SAM2’s original ability to segment and track objects in natural scenes.
Table 3: Comparisons between our method and existing approaches on CAD.
EGNet2019 [zhao2019egnet] ResNet-50 111.7 - 61.9 29.8 4.4 35.0 66.6 32.4 24.3
BASNet2019 [qin2019basnet] ResNet-50 87.1 - 63.9 34.9 5.4 39.4 77.3 39.3 29.3
CPD2019 [wu2019cascaded] ResNet-50 47.9 - 62.2 28.9 4.9 35.7 66.7 33.0 23.9
PraNet2020 [fan2020pranet] ResNet-50 32.6 - 62.9 35.2 4.2 39.7 76.3 37.8 29.0
SINet2020 [fan2020camouflaged] ResNet-50 48.9 - 63.6 34.6 4.1 39.5 77.5 38.1 28.3
SINet-V22021 [fan2021concealed] Res2Net-50 27.0 - 65.3 38.2 3.9 43.2 76.2 41.3 31.8
PNS-Net2021 [ji2021progressively] ResNet-50 142.9 - 65.5 32.5 4.8 41.7 67.3 38.4 29.0
RCRNet2019 [yan2019semi] ResNet-50 53.8 - 62.7 28.7 4.8 32.8 66.6 30.9 22.9
MG2021 [yang2021self] VGG 4.8 - 59.4 33.6 5.9 37.5 69.2 36.8 26.8
SLT-Net-LT2022 [cheng2022implicit] PVTv2-B5 82.3 - 69.6 48.1 3.0 52.4 84.5 49.3 40.2
ZoomNeXt2024 [pang2024zoomnext] PVTv2-B5 84.8 - 75.7 59.3 2.0 63.1 86.5 59.9 51.0
SAM22024 [ravi2025sam] Hiera-T 38.9 1-click 75.7 58.3 3.3 62.2 81.4 59.2 48.9
CamSAM2 Hiera-T 39.4 1-click 77.1 62.2 3.2 68.1 83.9 62.6 50.7
SAM22024 [ravi2025sam] Hiera-T 38.9 box 85.4 77.3 1.7 79.5 95.1 77.8 66.7
CamSAM2 Hiera-T 39.4 box 87.2 79.5 1.3 81.4 96.3 79.6 69.2
Results on CAD.
We evaluate the zero-shot performance of CamSAM2 and SAM2 on the CAD using Hiera-T backbone with point and box prompts, as shown in Tab.˜3. CamSAM2 consistently outperforms SAM2, especially with the 1-click prompt, achieving improvements of 3.4 mDice and 1.8 mIoU. With the box prompt, it also shows clear gains of 1.8 mDice and 2.5 mIoU. These results highlight CamSAM2’s superior zero-shot performance, demonstrating its effectiveness for segmentation tasks with minimal user input.
Results on SUN-SEG.
As shown in Tab.˜5, CamSAM2 consistently outperforms SAM2 across all metrics on both SUN-SEG-Easy and SUN-SEG-Hard. Notably, it improves mDice by 10.7 on SUN-SEG-Easy (from 73.6 to 84.3) and by 19.6 on SUN-SEG-Hard (from 61.0 to 80.6), demonstrating strong capability in segmenting camouflaged polyps. These results confirm that CamSAM2 significantly enhances SAM2 across varying tasks of camouflage, highlighting its effectiveness and generalizability in VCOS.
SUN-SEG-Easy
SAM2 [ravi2025sam] 83.4 71.6 83.0 73.6
CamSAM2 88.3 82.6 93.4 84.3
SUN-SEG-Hard
SAM2 [ravi2025sam] 75.5 58.4 73.4 61.0
CamSAM2 86.4 78.2 91.2 80.6
Decam.
Token
IOF EOF OPG mDice $\uparrow$ mIoU $\uparrow$
52.1 44.8
[…]
55.2 47.5
[…]
64.3 54.6
4.3 Ablation studies
To understand the impact of each component in CamSAM2, we conduct ablation studies on MoCA-Mask using the Hiera-T backbone with the 1-click prompt. The goal is to measure the contributions of key components, including the decamouflaged token, IOF, EOF, and OPG. Additionally, we evaluated the effects of different distance metrics and prototype numbers in the OPG process.
Impact of key components.
As shown in Tab.˜5, each main component in CamSAM2 clearly contributes to its high performance. Starting from baseline (SAM2), adding the decamouflaged token alone improves mDice from 52.1 to 54.9 and mIoU from 44.8 to 47.0. Adding IOF further raises mDice to 55.2 and mIoU to 47.5. Using EOF brings mDice to 55.9 and mIoU to 47.9. With all components included, the model performs the best, achieving 64.3 on mDice and 54.6 on mIoU.
     Distance Metric mDice $\uparrow$ mIoU $\uparrow$
Euclidean 61.9 52.7
Cosine 64.3 54.6
     Prototypes ($k$) mDice $\uparrow$ mIoU $\uparrow$
3 60.2 51.8
[…]
7 60.6 50.8
Effect of distance metric.
We compare different distance metrics for k-means clustering in OPG, as shown in Tab.˜7. Cosine distance performs better than Euclidean distance, likely due to its effectiveness in grouping camouflaged features by angular relationships rather than direct distances.
Influence of number of prototypes $k$.
We examine the impact of the number of prototypes $k$, as shown in Tab.˜7. The results show that both fewer or higher numbers of prototypes will reduce the performance due to under-representation or redundancy, respectively. It is observed that $k=5$ is found to be optimal for capturing essential informative details in camouflaged features.
5 Conclusion
Acknowledgments
This project was partially supported by grant #2022-279 of the Strategic Focus Area “Personalized Health and Related Technologies (PHRT)” of the ETH Domain (Swiss Federal Institutes of Technology).
Appendix A Appendix
A.1 Training details
During the training process, each training video clip consists of 8 frames, the input frames are resized to $1024\times 1024$, and the ground truths are resized to $256\times 256$ since the raw predicted logits are 1/4 of the original size. We train CamSAM2 with a batch size of 4. The predicted final mask is interpolated to $1024\times 1024$ and then encoded with the visual feature of the current frame as the memory feature in the memory bank for subsequent frames.
A.2 Comparative analysis across backbones and prompt types
Table 8: Detailed comparisons between CamSAM2 and SAM2 on MoCA-Mask across various backbones and prompt types. CamSAM2 consistently outperforms SAM2 for all cases. Improvements are highlighted in dark green.